an airplane is flying directly south at a speed of 195 m/s when a strong 65 m/s gust of wind heading due west affects the flight of the plane. what is the resulting speed of the airplance? and the resulting direction of the airplane?
1
Expert's answer
2020-12-21T10:43:29-0500

"v=\sqrt{65^2+195^2}\approx206(m/s)"


"\tan\alpha=\frac{195}{65}=3\to \alpha=71.6°" (south of west)